# Investigating charmed hybrid baryons via QCD sum rules

### v1

@@ Abstract
In the above expressions, we have evaluated the two-point correlation function up to the twelfth dimension, including the quark condensate $\langle\bar{q}q\rangle$, the two-gluon condensate $\langle g_{s}^{2}GG\rangle$, the three-gluon condensate $\langle g_{s}^{3}G^{3}\rangle$, the quark-gluon mixed condensate $\langle g_{s}\bar{q}\sigma Gq\rangle$, and their combinations.
Additionally, we have evaluated the sum rule result at the $\mathcal{O}(1/m_{c})$ order, and the mass of the $\Sigma_{cg}^{+}$ state is given by
$M_{+}(\omega_{c},T)=m_{c}+\bar{\Lambda}_{+}(\omega_{c},T)+\delta M_{+}(\omega_{c},T)\,.$ (13)
There are two free parameters, the threshold value $\omega_{c}$ and the Borel mass $T$, and we need to find their proper working regions. Detailed analyses are provided in the supplemental material, and the numerical results for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are provided in Table 1.
Table 1: Parameters of the charmed hybrid baryons belonging to the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$, calculated using the QCD sum rule method within the framework of heavy quark effective theory.
(GeV) (GeV) (GeV) (GeV) (GeV) (GeV4)
$[\bar{3}_{f},\bar{3}_{c},0,1,G]$ 3.20 $0.400\leq T\leq 0.410$ $2.72^{+0.17}_{-0.15}$ $\Lambda_{cg}(1/2^{+})$ $4.28^{+0.21}_{-0.19}$ $0.31^{+0.06}_{-0.05}$ $0.155^{+0.031}_{-0.024}$
$\Lambda_{cg}(3/2^{+})$ $4.59^{+0.24}_{-0.20}$ $0.045^{+0.009}_{-0.007}$
3.45 $0.409\leq T\leq 0.437$ $2.76^{+0.11}_{-0.09}$ $\Xi_{cg}(1/2^{+})$ $4.38^{+0.14}_{-0.12}$ $0.27^{+0.05}_{-0.04}$ $0.176^{+0.023}_{-0.019}$
$\Xi_{cg}(3/2^{+})$ $4.65^{+0.15}_{-0.13}$ $0.051^{+0.007}_{-0.005}$
$[6_{f},\bar{3}_{c},1,0,G]$ $2.80$ $0.394\leq T\leq 0.401$ $2.25\pm 0.11$ $\Sigma_{cg}(1/2^{+})$ $4.06^{+0.16}_{-0.15}$ $0.093^{+0.014}_{-0.012}$
3.05 $0.373\leq T\leq 0.425$ $2.32^{+0.14}_{-0.06}$ $\Xi^{\prime}_{cg}(1/2^{+})$ $4.17^{+0.16}_{-0.10}$ $0.109^{+0.019}_{-0.011}$
3.30 $0.357\leq T\leq 0.447$ $2.44^{+0.15}_{-0.08}$ $\Omega_{cg}(1/2^{+})$ $4.34^{+0.17}_{-0.11}$ $0.191^{+0.041}_{-0.020}$
$[6_{f},\bar{3}_{c},1,1,G]$ 2.95 $0.368\leq T\leq 0.398$ $2.46^{+0.30}_{-0.14}$ $\Sigma_{cg}(1/2^{+})$ $3.46^{+0.33}_{-0.21}$ $1.07^{+0.17}_{-0.14}$ $0.079^{+0.033}_{-0.030}$
$\Sigma_{cg}(3/2^{+})$ $4.54^{+0.38}_{-0.20}$ $0.046^{+0.019}_{-0.018}$
3.20 $0.351\leq T\leq 0.423$ $2.50^{+0.45}_{-0.08}$ $\Xi^{\prime}_{cg}(1/2^{+})$ $3.60^{+0.48}_{-0.16}$ $1.07^{+0.19}_{-0.13}$ $0.090^{+0.073}_{-0.010}$
$\Xi^{\prime}_{cg}(3/2^{+})$ $4.67^{+0.57}_{-0.12}$ $0.052^{+0.042}_{-0.006}$
3.45 $0.351\leq T\leq 0.447$ $2.62^{+0.28}_{-0.08}$ $\Omega_{cg}(1/2^{+})$ $3.77^{+0.31}_{-0.16}$ $1.07^{+0.16}_{-0.13}$ $0.158^{+0.072}_{-0.016}$
$\Omega_{cg}(3/2^{+})$ $4.84^{+0.35}_{-0.12}$ $0.091^{+0.042}_{-0.009}$
$[6_{f},\bar{3}_{c},1,2,G]$ 3.50 $0.418\leq T\leq 0.429$ $2.98^{+0.16}_{-0.14}$ $\Sigma_{cg}(3/2^{+})$ $5.36^{+0.24}_{-0.20}$ $-0.70^{+0.11}_{-0.11}$ $0.364^{+0.063}_{-0.051}$
$\Sigma_{cg}(5/2^{+})$ $4.66^{+0.21}_{-0.18}$ $0.123^{+0.021}_{-0.017}$
3.75 $0.428\leq T\leq 0.457$ $3.04^{+0.10}_{-0.09}$ $\Xi^{\prime}_{cg}(3/2^{+})$ $5.45^{+0.16}_{-0.14}$ $-0.67^{+0.09}_{-0.10}$ $0.422^{+0.052}_{-0.043}$
$\Xi^{\prime}_{cg}(5/2^{+})$ $4.78^{+0.15}_{-0.13}$ $0.143^{+0.018}_{-0.015}$
4.00 $0.434\leq T\leq 0.484$ $3.16\pm 0.09$ $\Omega_{cg}(3/2^{+})$ $5.63^{+0.15}_{-0.14}$ $-0.65^{+0.09}_{-0.10}$ $0.740^{+0.080}_{-0.073}$
$\Omega_{cg}(5/2^{+})$ $4.98^{+0.14}_{-0.13}$ $0.250^{+0.027}_{-0.025}$
Summary and discussions — In recent years, substantial theoretical and experimental efforts have been devoted to the study of glueballs and hybrid mesons. However, their internal structures remain elusive, underscoring the necessity for continued investigations—particularly into the comparatively less explored hybrid baryons. In this work, we employ the QCD sum rule method within the framework of heavy quark effective theory to investigate charmed hybrid baryons with various quantum numbers. We systematically construct twenty-eight interpolating currents for these baryons and utilize seven of them to compute the masses of nineteen states with quark-gluon configurations $qqcg$, $qscg$, and $sscg$ ($q=u/d$). The resulting mass predictions are summarized in Table 1.
The charmed hybrid baryons belonging to the doublet $[6_{f},\bar{3}c,1,1,G]$ comprise six states: $\Sigma_{cg}(1/2^{+})$, $\Sigma_{cg}(3/2^{+})$, $\Xi^{\prime}_{cg}(1/2^{+})$, $\Xi^{\prime}_{cg}(3/2^{+})$, $\Omega_{cg}(1/2^{+})$, and $\Omega_{cg}(3/2^{+})$. Among these, the three spin-$1/2$ states are predicted to be the lowest in mass:
$M_{\Sigma_{cg}(1/2^{+})}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$
$M_{\Xi^{\prime}_{cg}(1/2^{+})}$ $=$ $3.60^{+0.48}_{-0.16}~{}\rm{GeV}\,,$
$M_{\Omega_{cg}(1/2^{+})}$ $=$ $3.77^{+0.31}_{-0.16}~{}\rm{GeV}\,.$
These hybrid baryons can decay via the excitation of a color-octet $\bar{q}q$ or $\bar{s}s$ pair from the valence gluon, followed by the recombination of the remaining color-octet clusters—$qqc$, $qsc$, or $ssc$—with the excited $\bar{q}q$ or $\bar{s}s$ pair into color-singlet baryons and mesons. Representative decay channels for these states are listed in Table 2, offering viable avenues for their experimental identification in future studies.
Table 2: Possible decay channels of the charmed hybrid baryons belonging to the doublet $[6_{f},\bar{3}_{c},1,1,G]$.
States Decay patterns
$\Sigma_{cg}(1/2^{+})$ $ND$, $\Sigma D_{s}$, $ND^{*}$, $\Sigma D_{s}^{*}$, $\Delta D$, $\Sigma^{*}D_{s}$
$\Omega_{cg}(1/2^{+})$ $\Xi D$, $\Xi D^{*}$, $\Xi^{*}D$, $\Omega D_{s}$, $\Xi^{*}D^{*}$
ACKNOWLEDGMENTS
Supplemental Material
This supplemental material provides additional details on our framework. In this Letter we apply the QCD sum rule method to study the charmed hybrid baryons within the framework of heavy quark effective theory. Their interpolating currents can generally be constructed by using one charm quark field $h_{v}^{a}(x)$, two light quark fields $q_{a}(x)$ and $q_{b}(x)$, and one gluon field strength tensor $G_{\mu\nu}^{n}(x)$ (or one dual gluon field strength tensor $\widetilde{G}_{\mu\nu}^{n}(x)=G^{n,\rho\sigma}(x)\times\epsilon_{\mu\nu\rho\sigma}/2$), with the color indices $a/b=1\cdots 3$ and $n=1\cdots 8$ as well as the Lorentz indices $\mu/\nu=0\cdots 3$. The (dual) gluon field is a color octet, so the combination of the three quark fields should also form a color octet, with two possible combinations:
As summarized in Fig. 1, we identify a total of twenty-eight charmed hybrid baryon currents, denoted as $J^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}/\widetilde{J}^{\alpha_{1}\cdots\alpha_{J-1/2}}_{F,C,j_{qq},j_{l},J^{P}}$, which can be categorized into eight multiplets, denoted as $[F,C,j_{qq},j_{l},G/\widetilde{G}]$. The explicit expressions of these currents are given as follows:
  • The doublet $[\bar{3}_{f},\bar{3}_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[\bar{3}_{f},\bar{3}_{c},0,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
  • The singlet $[6_{f},\bar{3}_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $[6_{f},\bar{3}_{c},1,0,\widetilde{G}]$ contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
  • The doublet $[6_{f},\bar{3}_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[6_{f},\bar{3}_{c},1,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    ×$~{}~{}~{}~{}\times~{}~{}(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}G_{\mu\nu}^{n}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}\times~{}~{}(g^{\alpha\nu}_{t}-{1\over 3}\gamma_{t}^{\alpha}\gamma_{t}^{\nu})g_{s}{\lambda_{ec}^{n}\over 2}\widetilde{G}_{\mu\nu}^{n}h_{v}^{c}\,.$
  • The doublet $[6_{f},\bar{3}_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $[6_{f},\bar{3}_{c},1,2,\widetilde{G}]$ contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
  • The doublet $[6_{f},6_{c},0,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[6_{f},6_{c},0,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\sigma_{t}^{\mu\nu}g_{s}{\lambda_{eb}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,,$
  • The singlet $[\bar{3}_{f},6_{c},1,0,G]$ contains one state with $J^{P}=1/2^{+}$, and its dual singlet $[\bar{3}_{f},6_{c},1,0,\widetilde{G}]$ contains another state with $J^{P}=1/2^{-}$. Their corresponding currents are
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,.$
  • The doublet $[\bar{3}_{f},6_{c},1,1,G]$ contains two states with $J^{P}=1/2^{+}$ and $3/2^{+}$. Its dual doublet $[\bar{3}_{f},6_{c},1,1,\widetilde{G}]$ contains two other states with $J^{P}=1/2^{-}$ and $3/2^{-}$. Their corresponding currents are
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\mu\nu}^{m}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\times~{}~{}\gamma_{t}^{\nu}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\mu\nu}^{m}h_{v}^{c}\,,$
  • The doublet $[\bar{3}_{f},6_{c},1,2,G]$ contains two states with $J^{P}=3/2^{+}$ and $5/2^{+}$. Its dual doublet $[\bar{3}_{f},6_{c},1,2,\widetilde{G}]$ contains two other states with $J^{P}=3/2^{-}$ and $5/2^{-}$. Their corresponding currents are
    ×$~{}\times~{}~{}\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    ×$~{}\times~{}~{}\Gamma_{J=2}^{\alpha\mu_{4},\mu_{1}\nu_{2}}\gamma_{t}^{\mu_{4}}\gamma_{t}^{\nu_{1}}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}\times~{}~{}\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}G_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,,$
    ×$~{}~{}~{}~{}\times~{}~{}\Gamma_{J={5\over 2}}^{\alpha\beta,\mu_{1}\nu_{2}}\gamma_{t}^{\nu_{1}}\gamma_{5}g_{s}{\lambda_{ec}^{m}\over 2}\widetilde{G}_{\nu_{1}\nu_{2}}^{m}h_{v}^{c}\,.$
In the above expressions, $\gamma_{t}^{\mu}=\gamma^{\mu}-v\!\!\!/v^{\mu}$, $g_{t}^{\mu\nu}=g^{\mu\nu}-v^{\mu}v^{\nu}$, $\sigma_{t}^{\mu\nu}=\sigma^{\mu\nu}-\sigma^{\mu\alpha}v_{\alpha}v^{\nu}-\sigma^{\alpha\nu}v_{\alpha}v^{\mu}$, and $G_{\mu\nu}^{n}=\partial_{\mu}A_{\nu}^{n}-\partial_{\nu}A_{\mu}^{n}+g_{s}f^{npq}A_{p,\mu}A_{q,\nu}$; $\Gamma_{J=2}^{\alpha\beta,\mu\nu}$ and $\Gamma_{J=5/2}^{\alpha\beta,\mu\nu}$ are the $J=2$ and $J=5/2$ projection operators:
In this Letter we apply the QCD sum rule method to systematically study the charmed hybrid baryon currents listed above within the framework of heavy quark effective theory. In particular, the numerical results for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are presented in Table 1. As discussed previously, these currents can couple to charmed hybrid baryons with both positive and negative parities. For example, the positive-parity current $J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}$ generally couples to both the positive-parity state $B_{+}$ and the negative-parity state $B_{-}$ through
$\langle 0|J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}|B_{+}\rangle$ $=$ $f^{+}_{F,C,j_{qq},j_{l},J^{+}}u^{\alpha_{1}\cdots\alpha_{J-1/2}}\,,$ (46)
$\langle 0|J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}|B_{-}\rangle$ $=$ $f^{-}_{F,C,j_{qq},j_{l},J^{+}}\gamma_{5}u^{\alpha_{1}\cdots\alpha_{J-1/2}}\,.$ (47)
Its two-point correlation function can generally be written as
$\Pi_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2},\beta_{1},\cdots\beta_{J-1/2}}(\omega)$ (48)
$=$ $i\int d^{4}xe^{ikx}\langle 0|T[J_{F,C,j_{qq},j_{l},J^{+}}^{\alpha_{1}\cdots\alpha_{J-1/2}}(x)\bar{J}_{F,C,j_{qq},j_{l},J^{+}}^{\beta_{1},\cdots\beta_{J-1/2}}(0)]|0\rangle$
@@ ACKNOWLEDGMENTS
$\times$ $\Big{(}{1+v\!\!\!/\over 2}\Pi^{+}_{F,C,j_{qq},j_{l},J^{+}}(\omega)+{1-v\!\!\!/\over 2}\Pi^{-}_{F,C,j_{qq},j_{l},J^{+}}(\omega)\Big{)}\,,$
where $\mathbb{S}[\cdots]$ denotes the symmetrization and the subtraction of trace terms in the sets ($\alpha_{1}\cdots\alpha_{J-1/2}$) and ($\beta_{1}\cdots\beta_{-1/2}$).
At the quark-gluon level, we calculate Eq. (48) using the operator product expansion (OPE) method, and the sum rule equations for the multiplets $[\bar{3}_{f},\bar{3}_{c},0,1,G]$, $[6_{f},\bar{3}_{c},1,0,G]$, $[6_{f},\bar{3}_{c},1,1,G]$, and $[6_{f},\bar{3}_{c},1,2,G]$ are provided in the supplemental Mathematica file “OPE.nb”. During the calculations, we evaluate the OPE at the leading order, and the sum rule equation for the current $J_{\Sigma_{cg},\bar{3}_{c},1,1,1/2^{+}}$ is given in Eq. (10); we also evaluate the OPE at the $\mathcal{O}(1/m_{c})$ order, and the sum rule equations for this current are
@@
Note that in the above expressions, we have considered only the $\Sigma_{cg}^{+}$ state with the positive parity and neglected the $\Sigma_{cg}^{-}$ state with the negative parity.
To calculate Eq. (50) and Eq. (50), we consider the following Lagrangian from the heavy quark effective theory
$\mathcal{K}$ $=$ $\bar{h}_{v}^{c}(iD_{t})^{2}h_{v}^{c}\,,$ (52)
$\times$ $\left({1+v\!\!\!/\over 2}\delta_{\mathcal{O}}\Pi^{+}_{F,C,j_{qq},j_{l},J^{P}}+{1-v\!\!\!/\over 2}\delta_{\mathcal{O}}\Pi^{-}_{F,C,j_{qq},j_{l},J^{P}}\right)\,,$
Based on Eq. (51), we can express Eq. (55) at the hadron level as
where $K_{+}$, $\Sigma_{+}$, and $d_{M}$ are defined as
$d_{M}$$\equiv$ $d_{j,j_{l}}\,,$ (58)
$d_{j_{l}-1/2,j_{l}}$ $=$ $2j_{l}+2\,,$
$d_{j_{l}+1/2,j_{l}}$ $=$ $-2j_{l}\,.$
After performing the double Borel transformation to convert $\omega$ and $\omega^{\prime}$ into $T_{1}$ and $T_{2}$, and setting these two Borel parameters equal, we obtain Eq. (50) and Eq. (50). These two equations represent the sum rule results at the $\mathcal{O}(1/m_{c})$ order, from which we derive
$\delta M_{+}(\omega_{c},T)=-\frac{1}{4m_{c}}\left(K_{+}(\omega_{c},T)+d_{M}C_{mag}\Sigma_{+}(\omega_{c},T)\right)\,.$ (59)
$\langle\bar{q}q\rangle$ $=$ $-(0.240\pm 0.010)^{3}~{}\rm{GeV}\,,$
$\langle\bar{s}s\rangle$ $=$ $(0.8\pm 0.1)\times\langle\bar{q}q\rangle\,,$
$\alpha_{s}(\mu=M_{\tau})$ $=$ $0.33\,,$
$\overline{m}_{c}(\mu=m_{c})$ $=$ ${1.23\pm 0.09}~{}\rm{GeV}\,,$ (61)
$m_{c}(\mu=2~{}{\rm GeV})$ $=$ $\overline{m}_{c}({\alpha_{s}(\mu)\over\alpha_{s}(\overline{m}_{c})})^{12/25}=1.10\,,$
$m_{s}(\mu=2~{}{\rm GeV})$ $=$ $93^{+11}_{-~{}5}~{}\rm{MeV}\,.$
The sum rule equations listed in Eq. (10), Eq. (50), and Eq. (50) contain two free parameters: the threshold value $\omega_{c}$ and the Borel mass $T$. Their appropriate working regions are determined based on three criteria: a) sufficient convergence of the OPE, b) a sufficiently large pole contribution, and c) weak dependence of the sum rule results on these parameters.
Figure 2: CVG(′,′′) and PC as functions of the Borel mass $T$, with the threshold value set to be $\omega_{c}=2.95$ GeV. These curves are derived using the current $J_{\Sigma_{cg}^{+},\bar{3}_{c},1,1,1/2^{+}}$, which belongs to the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
$\bar{\Lambda}_{+}$ $=$ $2.46^{+0.30}_{-0.14}~{}\rm{GeV}\,,$ (64)
$f_{+}$ $=$ $0.079^{+0.033}_{-0.030}~{}\rm{GeV}^{4}\,.$ (65)
$K_{+}$ $=$ $-2.72^{+0.19}_{-0.26}~{}\rm{GeV}^{2}\,,$ (66)
$\Sigma_{+}$ $=$ $2.86^{+0.31}_{-0.25}~{}\rm{GeV}^{2}\,.$ (67)
Their variations are shown in Fig. 3 as functions of the Borel mass $T$, where the dependence is weak and remains within an acceptable range inside the Borel window $0.37~{}\rm{GeV}\leq T\leq 0.40~{}\rm{GeV}$. From this analysis, we extract the mass of the $\Sigma_{cg}^{+}$ state as
$M_{+}$ $=$ $3.46^{+0.33}_{-0.21}~{}\rm{GeV}\,,$ (68)
@@
Figure 3: Variations of $\bar{\Lambda}_{+}$, $K_{+}$, and $\Sigma_{+}$ as functions of the Borel mass $T$, where the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~{}\mathrm{GeV}$, $2.95~{}\mathrm{GeV}$, and $3.05~{}\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $[\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.

### v2

@@ III Results at the leading order
In the above expressions, we have evaluated the two-point correlation function up to the twelfth dimension, including the quark condensate $\langle\bar{q}q\rangle$, the two-gluon condensate $\langle g_{s}^{2}GG\rangle$, the three-gluon condensate $\langle g_{s}^{3}G^{3}\rangle$, the quark-gluon mixed condensate $\langle g_{s}\bar{q}\sigma Gq\rangle$, and their combinations. The sum rule equations for other currents belonging to the multiplets $\bar{3}_{f},\bar{3}_{c},0,1,G]$, $6_{f},\bar{3}_{c},1,0,G]$, $6_{f},\bar{3}_{c},1,1,G]$, and $6_{f},\bar{3}_{c},1,2,G]$ are provided in the supplemental Mathematica file “OPE.nb”.
$\langle\bar{q}q\rangle$ $=$ $-(0.240\pm 0.010)^{3}~\rm{GeV}\,$,
$\langle\bar{s}s\rangle$ $=$ $(0.8\pm 0.1)\times\langle\bar{q}q\rangle\,$,
$\alpha_{s}(\mu=M_{\tau})$ $=$ $0.33\,$,
$\overline{m}_{c}(\mu=m_{c})$ $=$ ${1.23\pm 0.09}~\rm{GeV}\,$, (46)
$m_{c}(\mu=2~{\rm GeV})$ $=$ $\overline{m}_{c}({\alpha_{s}(\mu)\over\alpha_{s}(\overline{m}_{c})})^{12/25}=1.10~{\rm GeV}\,$,
$m_{s}(\mu=2~{\rm GeV})$ $=$ $93^{+11}_{-~5}~\rm{MeV}\,.$
The sum rule equation listed in Eq. (42) contains two free parameters: the threshold value $\omega_{c}$ and the Borel mass $T$. Their appropriate working regions are determined based on three criteria: a) sufficient convergence of the OPE, b) a sufficiently large pole contribution, and c) weak dependence of the sum rule results on these parameters.
Figure 2: CVG(′,′′) and PC as functions of the Borel mass $T$, with the threshold value set to be $\omega_{c}=2.95$ GeV. These curves are derived using the current $J_{\Sigma_{cg}^{+},\bar{3}_{c},1,1,1/2^{+}}$, which belongs to the doublet $\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
$\bar{\Lambda}_{+}$ $=$ $2.46^{+0.15}_{-0.14}~\rm{GeV}\,$, (49)
$f_{+}$ $=$ $0.079^{+0.015}_{-0.013}~\rm{GeV}^{4}\,.$ (50)
Their variations are shown in Fig. 3 as functions of the Borel mass $T$, where the dependence is weak and remains within an acceptable range inside the Borel window $0.37~\rm{GeV}\leq T\leq 0.40~\rm{GeV}$.
Figure 3: Variations of (a) $\bar{\Lambda}_{+}$ and (b) $f_{+}$ as functions of the Borel mass $T$, where the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~\mathrm{GeV}$, $2.95~\mathrm{GeV}$, and $3.05~\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.
IV Results at the ${\mathcal{O}}(1/m_{Q})$ order
In this section we extend the analysis to include ${\mathcal{O}}(1/m_{Q})$ corrections by considering the following Lagrangian from the heavy quark effective theory
$\mathcal{K}$ $=$ $\bar{h}_{v}^{c}(iD_{t})^{2}h_{v}^{c}\,$, (52)
@@ IV Results at the ${\mathcal{O}}(1/m_{Q})$ order
$\times$ $\left({1+v\!\!\!/\penalty 50\over 2}\delta_{\mathcal{O}}\Pi^{+}_{F,C,j_{qq},j_{l},J^{P}}+{1-v\!\!\!/\penalty 50\over 2}\delta_{\mathcal{O}}\Pi^{-}_{F,C,j_{qq},j_{l},J^{P}}\right)\,$,
Based on Eq. (51), we can express Eq. (55) at the hadron level as
where $K_{+}$, $\Sigma_{+}$, and $d_{M}$ are defined as
$d_{M}$ $\equiv$ $d_{j,j_{l}}\,$, (58)
$d_{j_{l}-1/2,j_{l}}$ $=$ $2j_{l}+2\,$,
$d_{j_{l}+1/2,j_{l}}$ $=$ $-2j_{l}\,.$
The correlation functions in Eq. (55) can also be evaluated at the quark-gluon level using the operator product expansion. After performing the double Borel transformation to convert $\omega$ and $\omega^{\prime}$ into $T_{1}$ and $T_{2}$, and setting these two Borel parameters equal, we arrive at
@@
The Borel mass dependence of these sum rule equations is shown in Fig. 4. Both $K$ and $\Sigma$ exhibit mild variation with $T$ in the working region $2.7~\rm{GeV}\leq\omega_{c}\leq 3.2~\rm{GeV}$. From this region, we extract the following numerical results:
$K_{+}$ $=$ $-2.72^{+0.19}_{-0.20}~\rm{GeV}^{2}\,$, (61)
$\Sigma_{+}$ $=$ $2.87\pm 0.25~\rm{GeV}^{2}\,.$ (62)
These two values represent the sum rule results at the $\mathcal{O}(1/m_{c})$ order, from which we derive
$\delta M_{+}$ $=$ $-\frac{1}{2m_{c}}\left(K_{+}+d_{M}C_{mag}\Sigma_{+}\right)\,$, (63)
and further extract the mass of the $\Sigma_{cg}^{+}$ state as
$M_{\Sigma_{cg}(1/2^{+})}$ $\equiv$ $M_{+}=m_{c}+\overline{\Lambda}_{+}+\delta M_{+}$
$=$ $1.10^{+0.08}_{-0.08}~\mathrm{GeV}+2.46^{+0.15}_{-0.14}~\mathrm{GeV}-0.21^{+0.13}_{-0.12}~\mathrm{MeV}$
$=$ $3.36^{+0.27}_{-0.26}~\mathrm{GeV}\,$,
@@
In addition, the mass splitting between the $\Sigma_{cg}$ states of $J^{P}=1/2^{+}$ and $J^{P}=3/2^{+}$ is derived as
$\Delta M_{+}=M_{\Sigma_{cg}(3/2^{+})}-M_{\Sigma_{cg}(1/2^{+})}=2.16^{+0.32}_{-0.29}~\mathrm{GeV}\,.$ (65)
It is important to note that this mass splitting is rather large, an effect that should vanish in the heavy quark limit. As seen from Eq. (53), the term $\mathcal{S}$ is directly related to the mass splitting, and this term is significantly influenced by the gluon degrees of freedom. Therefore, the presence of a significant gluonic component in the charmed hybrid baryon might result in a substantial mass difference. Furthermore, as shown in Eq. (61) and Eq. (62), both the $K$ and $\Sigma$ terms are ${\mathcal{O}}(1/m_{Q})$ corrections, and their magnitudes are of the same order, indicating that the observed mass splitting is consistent with the ${\mathcal{O}}(1/m_{Q})$ correction scale. The effects of these gluonic contributions will be addressed further in future studies to quantify their impact more rigorously.
@@
Figure 4: Variations of (a) $K_{+}$ and (b) $\Sigma_{+}$ as functions of the Borel mass $T$, where the long-dashed, solid, and short-dashed curves correspond to fixed threshold values $\omega_{c}=2.85~\mathrm{GeV}$, $2.95~\mathrm{GeV}$, and $3.05~\mathrm{GeV}$, respectively. These results are obtained using the current $J_{\Sigma_{cg}^{+},\bar{3}c,1,1,1/2^{+}}$, which is associated with the doublet $\Sigma_{cg}^{+},\bar{3}_{c},1,1,G]$.